Calculate the components parallel (Fg//) and perpendicular( Fg ⊥) to the plane of a 150 N gravitational force acting on an object on a slope at 20° to the horizontal


1
Expert's answer
2021-02-01T08:06:54-0500
"F_{||}=W\sin{20}=150\sin{20}=51.3\ N\\
F_{⊥}=W\cos{20}=150\cos{20}=141.0\ N"